OpenStudy (anonymous):

The terms of a sequence {xn}∞ n=1 are not known exactly although it can be said that ∀n −n < xn < n. Prove that yn = xn/n^2 converges

OpenStudy (anonymous):

Hint: If \[-n <x_n<n\]Then dividing by n^2 yields:\[-\frac{1}{n} <\frac{x_n}{n^2} <\frac{1}{n}\]

OpenStudy (anonymous):

\[x+y+z /3\ge \sqrt[3]{xyz}\]